OpenStudy (anonymous):

HELP Gina, Sam and Robby all rented movies from the same video store. They each rented some dramas, comedies, and documentaries. Gina rented 11 movies total. Sam rented twice as many dramas, three times as many comedies, and twice as many documentaries as Gina. He rented 27 movies total. If Robby rented 19 movies total with the same number of dramas, twice as many comedies, and twice as many documentaries as Gina, how many movies of each type did Gina rent? Select one: a. 3 dramas, 5 comedies, and 3 documentaries b. 2 dramas, 6 comedies, and 3 documentaries c. 1 dramas, 4 comedies, and 6 d. 4 dramas, 3 comedies, and 4 documentaries

hero (hero):

c = comedies d = documentary r = dramas Gina's movies: c + d + r = 11 Sam's movies: 3c + 2d + 2r = 27 Robby's movies: 2c + 2d + r = 19

hero (hero):

Solve this system: c + d + r = 11 3c + 2d + 2r = 27 2c + 2d + r = 19

OpenStudy (anonymous):

how do you do that

hero (hero):

You can solve it using matrices or as a system of equations

OpenStudy (anonymous):

it would be D right?

hero (hero):

I only see a, b, and c as options

hero (hero):

But now that you mention it

hero (hero):

You can plug in the given values to see if they work for each equation

OpenStudy (anonymous):

so it is D

hero (hero):

I won't say which answer choice it is. Just plug in different numbers into the system. And be careful. Watch the variables. You have to make sure it works for each equation.

hero (hero):

This is multiple choice so I can't give you any hints. I helped you get this far. Now you have to figure out the rest on your own. The best way to proceed from here is to make sure you have it correct, then show your work to prove why the answer choice you chose was correct.